# Statistical Taylor Expansion111

### v1

@@ Abstract
Statistical Taylor expansion replaces the input precise variables in a conventional Taylor expansion with random variables each with known mean and deviation, to calculate the result mean and deviation. It is based on the uncorrelated uncertainty assumption [1]: Each input variable is measured independently with fine enough statistical precision, so that their uncertainties are independent of each other. Statistical Taylor expansion reviews that the intermediate analytic expressions can no longer be regarded as independent of each other, and the result of analytic expression should be path independent. This conclusion differs fundamentally from the conventional common approach in applied mathematics to find the best execution path for a result. This paper also presents an implementation of statistical Taylor expansion called variance arithmetic, and the tests on variance arithmetic.
@@ 1.2 Extension to Existing Statistics
  • The traditional variance-covariance framework [5] [6] only considers linear interaction between random variables through an analytic function. Statistical Taylor expansion generalizes it to higher order as Formula (2.19) in this paper.
For the discussion simplicity, and in line with most common statistical assumptions for error analysis [2][3], this paper assumes that each input uncertainty is Gaussian distributed, even though all formulas can be extended easily to generic distributions.
@@ 1.3.1 Conventional Floating-Point Arithmetic
Because a rounding error from lower digits quickly propagates to higher digits, the $10^{-7}$ precision of the 32-bit IEEE floating-point format [7][8][9] is usually not fine enough for calculations involving input data of $10^{-2}$ to $10^{-4}$ precision.
Self-censored rules are developed to avoid such rounding error propagation [13][14], such as avoiding subtracting results of large multiplication, as in Formula (1.2). However, these rules are not enforceable, and in many cases are difficult to follow, and even more difficult to quantify. So far, the study of rounding error propagation is concentrated on linear calculations [11] [12] [15], or on special cases [13] [16] [17], while the rounding errors are generally manifested as the lack of numerical stability [18].
The forward rounding error study [12] compares 1) the result with rounding error and 2) the ideal result without rounding error, such as comparing the result using 32-bit IEEE floating-point arithmetic and the corresponding result using 64-bit IEEE floating-point arithmetic [19]. The most recent such study gives an extremely optimistic view on numerical library errors as a fraction of the least significant bit of the significand of floating-point results [19]. However, such optimism does not seem to hold in the real statistical tests on the numerical library functions as presented in this paper.
@@
The conventional numerical approaches explore the path-dependency property of calculations based on conventional floating-point arithmetic, to find the best possible methods, such as in Gaussian elimination [12] [13] [14]. The implied statistical assumption is that the intermediate expressions are independent of each other. However, by showing that the intermediate expressions are dependent of each other, and the result should be path-independent, statistical Taylor expansion challenges this conventional wisdom. Variance arithmetic further provides a sole procedure to solve each analytic problem.
@@ 1.3.2 Interval Arithmetic
Interval arithmetic [14][20][21][22][23][24] is currently a standard method to track calculation uncertainty. Its goal is to ensure that the value x is absolutely bounded within its bounding range throughout the calculation. For example, Formula (1.3), (1.4), (1.5), and (1.6) gives the result of addition, subtraction, multiplication, and division, in which $\underline{x}$ and $\bar{x}$ are lower and upper bounds for $x$, respectively [24] [22].
$\left[x_{1}\right]+\left[x_{2}\right]=\left[\underline{x}_{1}+\underline{x}_{2},\bar{x}_{1}+\bar{x}_{2}\right];$ (1.3)
$\left[x_{1}\right]-\left[x_{2}\right]=\left[\underline{x}_{1}-\bar{x}_{2},\bar{x}_{1}-\underline{x}_{2}\right];$ (1.4)
$\left[x_{1}\right]\times\left[x_{2}\right]=\left[min(\underline{x}_{1}\underline{x}_{2},\underline{x}_{1}\bar{x}_{2},\bar{x}_{1}\underline{x}_{2},\bar{x}_{1}\bar{x}_{2}),max(\underline{x}_{1}\underline{x}_{2},\underline{x}_{1}\bar{x}_{2},\bar{x}_{1}\underline{x}_{2},\bar{x}_{1}\bar{x}_{2})\right];$ (1.5)
$0\notin\left[x_{2}\right]:\;\left[x_{1}\right]/\left[x_{2}\right]=\left[x_{1}\right]\times\left[1/\bar{x}_{2},1/\underline{x}_{2}\right];$ (1.6)
The bounding range used by interval arithmetic is not compatible with usual scientific and engineering measurements, which instead use the statistical mean and deviation to characterize uncertainty [2][3] 222 There is some attempt [25] to connect intervals in interval arithmetic to confidence interval or the equivalent so-called p-box in statistics. Because this attempt seems to rely heavily on 1) specific properties of the uncertainty distribution within the interval and/or 2) specific properties of the functions upon which the interval arithmetic is used, this attempt does not seem to be generic. If probability model is introduced to interval arithmetic to allow tiny bounding leakage, the bounding range is much less than the corresponding pure bounding range [15]. Anyway, these attempts seem to be outside the main course of interval arithmetic. .
Interval arithmetic only provides the worst case of uncertainty propagation. For instance, in addition, it gives the bounding range when the two input variables are +1 correlated [26]. However, if the two input variables are -1 correlated, the bounding range after addition reduces, which is the best case in random interval arithmetic [27] 333 The vast overestimation of bounding ranges in these two worst cases prompts the development of affine arithmetic [26][28], which traces error sources using a first-order model. Being expensive in execution and depending on approximate modeling even for such basic operations as multiplication and division, affine arithmetic has not been widely used. In another approach, random interval arithmetic [27] randomly chooses between the best-case and the worst-case intervals. Anyway, these attempts seem to be outside the main course of interval arithmetic. . Such worse assumption leads to gross overestimation of result uncertainties [1].
@@
Interval arithmetic lacks mechanism to reject a calculation, while a mathematical calculation always has a valid input range for a given targeted domain. For example, it forces branched results for $1/(x\pm\Delta x)$ or $\sqrt{x\pm\Delta x}$ when $0\in[x-\Delta x,x+\Delta x]$, while a context-sensitive uncertainty bearing arithmetic should reject such calculation naturally.
In contrast, variance arithmetic provides statistical bounding with a fixed bounding leakage of $5.73\times 10^{-7}$ probability. It has no dependency problem. Its statistical context rejects certain input intervals mathematically, such as inversion and square root of an interval containing zero.
@@ 1.3.3 Statistical Propagation of Uncertainty
If each input variable is a random variable, and the statistical correlation between the two input variables is known, statistical propagation of uncertainty gives the result mean and variance [29][30]. For example, Formula (1.7), (1.8), (1.9), and (1.10) gives the result of addition, subtraction, multiplication, and division, in which $\sigma$ is the deviation of a measured value $x$, $P$ is its precision, and $\gamma$ is the correlation between the two imprecise values:
$(x\pm\delta x)+(y\pm\delta y)$ $=(x+y)$$\pm\sqrt{\delta x^{2}+\delta y^{2}+2\delta x\delta y\gamma};$ (1.7)
$(x\pm\delta x)-(y\pm\delta y)$ $=(x-y)$$\pm\sqrt{\delta x^{2}+\delta y^{2}-2\delta x\delta y\gamma};$ (1.8)
$(x\pm\delta x)\times(y\pm\delta y)$ $=(x\times y)$$\pm|x\times y|\sqrt{{P_{x}}^{2}+{P(y)}^{2}+2P_{x}P(y)\gamma};$ (1.9)
$(x\pm\delta x)/(y\pm\delta y)$ $=(x/y)$$\pm|x/y|\sqrt{{P_{x}}^{2}+{P(y)}^{2}-2P_{x}P(y)\gamma};$ (1.10)
Statistical Taylor expansion has a different statistical assumption called uncorrelated uncertainty assumption [1]: Each input variable is measured independently with fine enough statistical precision, so that their uncertainties are independent of each other, even though the input variables could be highly correlated. This assumption is in line with most error analysis [2][3].
@@ 1.3.4 Significance Arithmetic
Significance arithmetic [31] tries to track reliable bits in an imprecise value during the calculation. In the two early attempts [32][33], the implementations of significance arithmetic are based on simple operating rules upon reliable bit counts, rather than on formal statistical approaches. They both treat the reliable bit counts as integers when applying their rules, while a reliable bit count could be a fractional number [34], so they both can cause artificial quantum reduction of significance. The significance arithmetic marketed by Mathematica [34] uses a linear error model that is consistent with a first-order approximation of interval arithmetic [14][22][23].
One problem of significance arithmetic is that itself can not properly specify the uncertainty [1]. For example, if the least significand bit of significand is used to specify uncertainty, then the representation has very coarse precision, such as $1\pm 10^{-3}=1024\times 2^{-10}$ [1]. Introducing limited bits calculated inside uncertainty cannot avoid this problem completely. Thus, the resolution of the conventional floating-point representation is desired. This is the reason why variance arithmetic has abandoned the significance arithmetic nature of its predecessor [1].
@@ 1.4 An Overview of This Paper
This paper presents the theory for statistical Taylor expansion, the implementation as variance arithmetic, and the tests of variance arithmetic. Section 1 compares statistical Taylor expansion and variance arithmetic with each known uncertainty-bearing arithmetic. Section 2 presents the theoretical foundation for statistical Taylor expansion Section 3 presents additional assumption for variance arithmetic. Section 4 discusses the standards and methods to validate variance arithmetic. Section 5 tests variance arithmetic for the most common math library functions. Section 6 applies variance arithmetic to adjugate matrix and matrix inversion. Section 7 applies variance arithmetic to process time-series data. Section 8 focuses on the effect of numerical library errors and shows that such library errors can be very significant. Section 9 applies variance arithmetic to regression. Section 10 provides a summary and discussion for variance arithmetic.
@@ 2.1 The Uncorrelated Uncertainty Assumption [1]
figure Allowed maximal correlation between two values vs. input precisions and independence standard (as shown in legend) for the independence uncertainty assumption of statistical Taylor expansion to be true. This figure has been published previously [1].
When there is a good estimation of the sources of uncertainty, the uncorrelated uncertainty assumption can be judged directly, e.g., if noise [2][3] is the major source of uncertainty, the uncorrelated uncertainty assumption is probably true. Otherwise, the uncorrelated uncertainty assumption can be judged by the correlation and the respectively precision of two measurements.
The correlated parts common to different measurements are regarded as signals, which can either be desired or unwanted. Let $X$, $Y$, and $Z$ denote three mutually independent random variables [5] with variance $V(X)$, $V(Y)$ and $V(Z)$, respectively. Let $\alpha$ denote a constant. Let $Cov()$ denote the covariance function. Let $\gamma$ denote the correlation between $(X+Y)$ and $(\alpha X+Z)$. And let:
@@
Formula (2.2) gives the correlation $\gamma$ between two random variables, each of which contains a completely uncorrelated part and a completely correlated part $X$, with $\eta$ being the average ratio between these two parts. Formula (2.2) can also be interpreted reversely: if two random variables are correlated by $\gamma$, each of them can be viewed hypothetically as containing a completely uncorrelated part and a completely correlated part, with $\eta$ being the average ratio between these two parts  [^4] The widely used common factor analysis is also based on this principle..
One special application of Formula (2.2) is the correlation between a measured signal and its true signal, in which noise is the uncorrelated part between the two. Figure 2.1 shows the effect of noise on the most significant two bits of a 4-bit measured signal when $\eta=1/4$. Its top chart shows a triangular waveform between 0 and 16 as a black line, and a white noise between -2 and +2, using the grey area. The measured signal is the sum of the triangle waveform and the noise. The middle chart of Figure 2.1 shows the values of the 3rd digit of the true signal as a black line, and the mean values of the 3rd bit of the measurement as a grey line. The 3rd bit is affected by the noise during its transition between 0 and 1. For example, when the signal is slightly below 8, only a small positive noise can turn the 3rd digit from 0 to 1. The bottom chart of Figure 2.1 shows the values of the 2nd digit of the signal and the measurement as a black line and a grey line, respectively. Figure 2.1 clearly shows that the correlation between the measurement and the true signal is less at the 2nd digit than at the 3rd digit. Quantitatively, according to Formula (2.2):
  1. 1.
@@
Figure 2.1 plots the relation of $\gamma$ vs. $P$ for each given $\gamma_{P}$ in Formula (2.4). When $\gamma_{P}$ is less than a predefined maximal threshold (e.g., 2%, 5% or 10%), the two values can be deemed uncorrelated to each other at the level of uncertainty. For each independence standard $\gamma_{P}$, there is a maximal allowed correlation between two values below which the uncorrelated uncertainty assumption of statistical Taylor expansion holds. The maximal allowed correlation is a function of the larger precision of the two values according to Formula (2.4). Figure 2.1 shows that for two precisely measured values, their correlation $\gamma$ is allowed to be quite high. Thus, the uncorrelated uncertainty assumption has much weaker statistical requirement than the assumption for independence arithmetic, which requires the two values to be independent of each other.
It is tempting to add noise to otherwise unqualified values to make them satisfy the uncertainties uncertainty assumption. As an extreme case of this approach, if two values are constructed by adding noise to the same signal, they are 50% correlated at the uncertainty level so that they will not satisfy the uncorrelated uncertainty assumption, which defines the upper bound for $\gamma_{P}$. 555 The $50\%$ curve in Figure 2.1 defines the maximal possible correlations between any two measured signals. This other conclusion of Formula (2.4) makes sense because the measurable correlation between two measurements should be limited by the precision of their measurements.
2.2 Addition and Subtraction
The Lyapunov form of the central limit theorem [5] states that if $X_{i}$ is a random variable with mean $\mu_{i}$ and variance $V_{i}$ for each $i$ among a series of $n$ mutually independent random variables, then with increased $n$, the sum $\sum\limits_{i}^{n}X_{i}$ converges in distribution to the Gaussian distribution with mean $\sum\limits_{i}^{n}\mu_{i}$ and variance $\sum\limits_{i}^{n}V_{i}$. In the context of the uncorrelated uncertainty assumption, uncertainty in general is expected to be Gaussian distributed [2] [5].
Even if rounding error is the sole source of uncertainty:
  • The rounding error of rounding to nearest is shown to be uniformly distributed between half of the least significant bit of an integer [1].
  • The central limit theorem converges the result distribution of addition and subtraction to a Gaussian distribution [1]. Such convergence to Gaussian distribution is very quick, e.g., after 4 rounds of addition and subtraction, the result distribution is already close to the corresponding Gaussian distribution with the same mean and deviation [1].
  • The application of center limit theorem to rounding error is not limited to addition and subtraction, because a multiplication operation contains multiple additions, while a division operation contains multiple subtractions [10]. Thus, rounding error is Gaussian distributed in practice.
For simplicity, define operator $\delta^{2}f(x)\equiv(\delta f(x))^{2}$. Formula (2.5) gives the result variance of addition and subtraction. Formula (2.6) gives the probability density function $\rho$ for $x\pm\delta x$ in general, in which $N(x)$ is the density function of the Normal distribution [5]. Formula (2.6) can be normalized as Formula (2.7).
$\delta^{2}(x\pm y)$ $=(\delta x)^{2}+(\delta y)^{2};$ (2.5)
$\rho(\tilde{x},x,\delta x)$ $=\frac{1}{\delta x}N(\frac{\tilde{x}-x}{\delta x});$ (2.6)
$\tilde{z}\equiv\frac{\tilde{x}-x}{\delta x}:\;\;\;$ $\rho(\tilde{x},x,\delta x)d\tilde{x}=N(\tilde{z})d\tilde{z};$ (2.7)
2.3 Multiplication
The result variance of multiplying $x\pm\delta x$ by a precise value $y$ is $y^{2}\delta^{2}x$ [5]. The result of multiplying $0\pm\delta x$ by $0\pm\delta y$ is a normal production distribution [5], which centers at 00 with variance $(\delta x)^{2}(\delta y)^{2}$. The general multiplication can be decomposed as Formula (2.8), which leads to Formula (2.9) [1].
$(x\pm\delta x)\times(y\pm\delta y)=(x+(0\pm\delta x))\times(y+(0\pm\delta y));$ (2.8)
$\delta^{2}(xy)$ $=x^{2}(\delta y)^{2}+y^{2}(\delta x)^{2}+(\delta x)^{2}(\delta y)^{2};$ (2.9)
Formula (2.9) is identical to Formula (1.9) for statistical propagation of uncertainty except their cross term, representing difference in their statistical requirements, respectively. Formula (2.9) is simpler and more elegant than Formula (1.5) for interval arithmetic multiplication, because the latter looks more like a wish. . The result of Formula (1.2) calculated by statistical Taylor expansion is $2\pm 2\sqrt{5}$. It is $2\pm 9$ for interval arithmetic.

### v2

@@ Abstract
Statistical Taylor expansion replaces the input precise variables in a conventional Taylor expansion with random variables each with known distribution, to calculate the result mean and deviation. It is based on the uncorrelated uncertainty assumption [1]: Each input variable is measured independently with fine enough statistical precision, so that their uncertainties are independent of each other. Statistical Taylor expansion reveals that the intermediate analytic expressions can no longer be regarded as independent of each other, and the result of analytic expression should be path independent. This conclusion differs fundamentally from the conventional common approach in applied mathematics to find the best execution path for a result. This paper also presents an implementation of statistical Taylor expansion called variance arithmetic, and the tests on variance arithmetic.
@@ 1.2 Extension to Existing Statistics
  • The traditional variance-covariance framework only considers linear interaction between random variables through an analytic function [5][6]. Statistical Taylor expansion generalizes it to higher order as Formula (2.3) in this paper.
@@ 1.3.1 Conventional Floating-Point Arithmetic
Because a rounding error from lower digits quickly propagates to higher digits, the $10^{-7}$ significance of the 32-bit IEEE floating-point format [7][8][9] is usually not fine enough for calculations involving input data of $10^{-2}$ to $10^{-4}$ precision. For complicated calculations, even the $10^{-16}$ significance of the 64-bit IEEE floating-point format [7][8][9] may not enough for inputs with $10^{-2}$ to $10^{-4}$ precision.
Self-censored rules are developed to avoid such rounding error propagation [13][14], such as avoiding subtracting results of large multiplication, as in Formula (1.2). However, these rules are not enforceable, and in many cases are difficult to follow, and even more difficult to quantify. So far, the study of rounding error propagation is concentrated on linear calculations [11][12][15], or on special cases [13][16][17], while the rounding errors are generally manifested as the mysterious and ubiquitous numerical instability [18].
The forward rounding error study [12] compares 1) the result with rounding error and 2) the ideal result without rounding error, such as comparing the result using 32-bit IEEE floating-point arithmetic and the corresponding result using 64-bit IEEE floating-point arithmetic [19]. The most recent such study gives an extremely optimistic view on numerical library errors as a fraction of the least significant bit of the significand of floating-point results [19]. However, such optimism does not seem to hold in the real statistical tests on the numerical library functions as presented in this paper. In contrast, the results of variance arithmetic agree well with all tests statistically.
@@
The conventional numerical approaches explore the path-dependency property of calculations based on conventional floating-point arithmetic, to find the best possible methods, such as in Gaussian elimination [12][13][14]. However, by showing that the analytic result of statistical input(s) should be path-independent, statistical Taylor expansion challenges this conventional wisdom.
@@ 1.3.2 Interval Arithmetic
Interval arithmetic [14][20][21][22][23][24] is currently a standard method to track calculation uncertainty. Its goal is to ensure that the value x is absolutely bounded within its bounding range throughout the calculation.
The bounding range used by interval arithmetic is not compatible with usual scientific and engineering measurements, which instead use the statistical mean and deviation to characterize uncertainty [2][3] 222 There is one attempt to connect intervals in interval arithmetic to confidence interval or the equivalent so-called p-box in statistics [25]. Because this attempt seems to rely heavily on 1) specific properties of the uncertainty distribution within the interval and/or 2) specific properties of the functions upon which the interval arithmetic is used, this attempt does not seem to be generic. If probability model is introduced to interval arithmetic to allow tiny bounding leakage, the bounding range is much less than the corresponding pure bounding range [15]. Anyway, these attempts seem to be outside the main course of interval arithmetic. .
Interval arithmetic only provides the worst case of uncertainty propagation. For instance, in addition, it gives the bounding range when the two input variables are +1 correlated [26]. However, if the two input variables are -1 correlated, the bounding range after addition reduces [27] 333 Such case is called the best case in random interval arithmetic. The vast overestimation of bounding ranges in these two worst cases prompts the development of affine arithmetic [26][28], which traces error sources using a first-order model. Being expensive in execution and depending on approximate modeling even for such basic operations as multiplication and division, affine arithmetic has not been widely used. In another approach, random interval arithmetic [27] randomly chooses between the best-case and the worst-case intervals, so that it can no longer guarantee bounding without leakage. Anyway, these attempts seem to be outside the main course of interval arithmetic. . Such worse assumption leads to overestimation of result uncertainties by order-of-magnitude [1].
@@
Interval arithmetic lacks mechanism to reject a calculation, while a mathematical calculation always has a valid input range. For example, it forces branched results for $1/(x\pm\Delta x)$ or $\sqrt{x\pm\Delta x}$ when $0\in[x-\Delta x,x+\Delta x]$, while a context-sensitive uncertainty bearing arithmetic should reject such calculation naturally.
In contrast, variance arithmetic provides statistical bounding range with a fixed bounding leakage of $5.73\times 10^{-7}$ probability. It has no dependency problem. Its statistical context rejects certain input intervals mathematically, such as inversion and square root of a statistical bounding range containing zero.
@@ 1.3.3 Statistical Propagation of Uncertainty
If each input variable is a random variable, and the statistical correlation between the two input variables is known, statistical propagation of uncertainty gives the result mean and variance [29][30].
Statistical Taylor expansion has a different statistical assumption called uncorrelated uncertainty assumption [1]: Each input variable is measured independently with fine enough precision, so that their uncertainties are independent of each other, even though the input variables could be highly correlated. This assumption is in line with most error analysis [2][3].
@@ 1.3.4 Significance Arithmetic
Significance arithmetic [31] tries to track reliable bits in an imprecise value during the calculation. In the two early attempts [32][33], the implementations of significance arithmetic are based on simple operating rules upon reliable bit counts, rather than on formal statistical approaches. They both hold the reliable a bit count as an integer when applying their rules, while a reliable bit count could be a fractional number [34], so they both can cause artificial quantum reduction of significance. The significance arithmetic marketed by Mathematica [34] uses a linear error model that is consistent with a first-order approximation of interval arithmetic [14][22][23].
One problem of significance arithmetic is that itself can not properly specify the uncertainty [1]. For example, if the least significand bit of significand is used to specify uncertainty, then the representation has very coarse precision, such as $1\pm 10^{-3}=1024\times 2^{-10}$ [1]. Introducing limited bits calculated inside uncertainty cannot avoid this problem completely. Thus, the resolution of the conventional floating-point representation is desired. This is why variance arithmetic has abandoned the significance arithmetic nature of its predecessor [1].
@@ 1.4 An Overview of This Paper
This paper presents the theory for statistical Taylor expansion, the implementation as variance arithmetic, and the tests of variance arithmetic. Section 1 compares statistical Taylor expansion and variance arithmetic with each known uncertainty-bearing arithmetic. Section 2 presents the theoretical foundation for statistical Taylor expansion. Section 3 presents variance arithmetic. Section 4 discusses the standards and methods to validate variance arithmetic. Section 5 tests variance arithmetic for the most common math library functions. Section 6 applies variance arithmetic to adjugate matrix and matrix inversion. Section 7 applies variance arithmetic to process time-series data. Section 8 focuses on the effect of numerical library errors and shows that such library errors can be significant. Section 9 applies variance arithmetic to regression. Section 10 provides a summary and discussion for variance arithmetic.
@@ 2.1 The Uncorrelated Uncertainty Assumption [1]
figure $\gamma$ vs $P$ for different $\gamma_{P}$ in Formula (2.4). If $\gamma_{P}$ is the maximal allowed correlation on the uncertainty level for uncorrelated uncertainty assumption to hold, then $\gamma$ is the maximal allowed correlation for two input to hold. For a given $\gamma_{P}$, $\gamma$ increases with finer $P$. This figure is a reproduction of Figure 2 in [1], with the P-axis in linear scale instead of in logarithmic scale, and with all texts improved.
The uncorrelated uncertainty assumption [1] states that the uncertainty of any input is statistically uncorrelated with any other input uncertainty. It is consistent with the common methods in processing experimental data [2][3], such as the common knowledge or belief that precision improves with the count $n$ as $1/\sqrt{n}$ during averaging.
However, the uncorrelated uncertainty assumption allows any two inputs to be highly correlated if each input is independently measured [1]:
Let $X$, $Y$, and $Z$ denote three mutually independent random variables [5] with variance $V(X)$, $V(Y)$ and $V(Z)$, respectively. Let $\alpha$ denote a constant. Let $Cov()$ denote the covariance function. Let $\gamma$ denote the correlation between $(X+Y)$ and $(\alpha X+Z)$. And let:
@@
Formula (2.2) gives the correlation $\gamma$ between two random variables, each of which contains a completely uncorrelated part and a completely correlated part $X$, with $\eta$ being the average ratio between these two parts. Formula (2.2) can also be interpreted reversely: if two random variables are correlated by $\gamma$, each of them can be viewed hypothetically as containing a completely uncorrelated part and a completely correlated part, with $\eta$ being the average ratio between these two parts  [^4] The widely used common factor analysis is also based on this principle.. The correlated parts common to different measurements are regarded as signals, which can either be desired or unwanted.
One special application of Formula (2.2) is the correlation between a measured signal and its true signal, in which noise is the uncorrelated part between the two. Figure 2.1 shows the effect of noise on the most significant two bits of a 4-bit measured signal when $\eta=1/4$. Its top chart shows a triangular waveform between 0 and 16 as a black line, and a white noise between -2 and +2, using the grey area. The measured signal is the sum of the triangle waveform and the noise. The middle chart of Figure 2.1 shows the values of the 3rd digit of the true signal as a black line, and the mean values of the 3rd bit of the measurement as a grey line. The 3rd bit is affected by the noise during its transition between 0 and 1. For example, when the signal is slightly below 8, only a small positive noise can turn the 3rd digit from 0 to 1. The bottom chart of Figure 2.1 shows the values of the 2nd digit of the signal and the measurement as a black line and a grey line, respectively. Figure 2.1 clearly shows that the correlation between the measurement and the true signal is less at the 2nd digit than at the 3rd digit. Quantitatively:
@@
Figure 2.1 plots the relation of $\gamma$ vs. $P$ for each given $\gamma_{P}$ in Formula (2.4). When $\gamma_{P}$ is less than a predefined maximal threshold (e.g., 2%, 5% or 10%), the two values can be deemed uncorrelated to each other at the level of uncertainty. For each independence standard $\gamma_{P}$, there is a maximal allowed correlation between two values below which the uncorrelated uncertainty assumption of statistical Taylor expansion holds. The maximal allowed correlation is a function of the larger precision of the two values according to Formula (2.4). Figure 2.1 shows that for two precisely measured values, their correlation $\gamma$ is allowed to be quite high.
In Formula (2.4), according to Formula (2.2), $P=\sqrt{P_{1}P_{2}}$, in which $P_{1}$ and $P_{2}$ are the precision of the two inputs, respectively. When $P=0$, $\gamma=1$, showing the implied statistical allowance when either or both inputs are precise.